I am to find a polynomial for the remaining area of a rectangular shaped corral measuring 23m by 17m, after a circular based silo of x meters is built inside of it. Thus far, I have A=23m X 17m - nr^2. Is this correct?
Answer by richwmiller(17219)

Is x the diameter of the silo or the radius?
if x is the radius then the silo takes up pi x^2
if x is the diameter then the silo takes up pi *(x/2)^2
We have 23*17=391 sq meters for the corral without the silo
A=391-pi * x^2
or
A=391-pi *(x/2)^2